User Tools

Site Tools


handover:aug029

aug029

General Notes*

  • 16:45 UT Kept losing network access while trying to fix Ke problems (live webpages would become unresponsive). Did not impact hb or yg erc for some reason. (Bryn)

Hobart 12m:

Disk VSN: HOB+0074

Data volume at beginning: 2694.897 GB

  • 18:00 UT Experiment started OK. Scan 265-1800, source 0454-234. (AK)

Katherine 12m:

Disk VSN: HOB+0059

Data volume at beginning: 0 GB

  • 18:07 UT experiment started late. Module had no vsn and internal light at ke didn't work, needed confirmation from on call person to give a place holder name. (AK)
  • 16:16 UT First scan 265-1816. Source 2255-282. (AK)
  • 16:01 UT DBBC pps_delay1 = 2048031248. ALARM: Delay through DBBC has changed. Monitor for stability and reset DBBC if it drifts. DBBC delay returned to regular levels after this one alarm, however, the autocorrelation spectra are awful. Reconfiguring DBBC. Missed scans 266-1600 onwards (Bryn)
  • 16:30 UT Reconfiguring DBBC did nothing autocorrelations still terrible (see attached), trying reboot of DBBC machine. (Bryn)
  • 18:00 UT Was unable to fix issues with DBBC, all scans from 266-1600 onwards bad.

Yarragadee 12m:

Disk VSN: HOB+0100

Data volume at beginning: 674.261 GB

  • 18:00 UT Experiment started OK. Scan 265-1800, source 0454-234. (AK)

Hobart 26m:

Disk VSN: HOB+1006

Data volume at beginning: 0 GB

  • 18:00 UT Lots of following errors
    • ERROR ch -369 IF3 lo is unlocked.
    • ERROR ch -371 IF3 alarm is on
    • ERROR ch -371 IF3 alarm is on
    • ERROR ch -309 v3 alarm is on
    • ERROR ch -309 v6 alarm is on
    • ERROR ch -309 v7 alarm is on
    • ERROR ch -309 vb alarm is on
    • ERROR ch -309 vc alarm is on
  • 1900 UT Jamie had a look and tried to fix it. He wasn't convinced that he had fixed the problem but the error messages disappeared. Brett has suggested power cycling the VC rack. Holding off on going to MT pleasant as there are currently no errors.
  • 20:11 UT disk_pos is 60 GB behind already. (AK)
  • 21:08 UT It appears the telescope didnt follow the schedule and stowed for some reason. First valid scan was 265-2108.
  • 00:05 UT Tried manually cycling power to VC rack, no luck - errors still persist. Jim L stowed telescope. (Bryn)
/home/www/auscope/opswiki/data/pages/handover/aug029.txt · Last modified: 2016/09/22 23:24 by Jesse Swan